what is the present value of quarterly cash flow (first cash flow comes in a quarter from today) of 1100$ (each) for 10 years? The APR is 6% compounded monthly.
Answer #1

Rate per quarter = (1+APR/12)12/4 -1 = (1 +6%/12)3 -1 = 1.5075%

Number of Periods = 4*10 = 40

PV of quarterly cash flow = PMT*(1-(1+r)-n)/r = 1100*(1-(1+1.5075%)-40)/1.5075% = 32,862.35

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well